The molecular mass of an organic compound carbon, hydrogen and nitrogen was found to be 59.on combustion, 0.13g of the compound gave 0.29g of CO2 and 0.18 of H2O. 0.145g of the compound on suitable treatment gave 22.4cm3 of N2.measured at s.t. p. Deduce the molecular formula of the compound.
1
Expert's answer
2020-02-19T07:57:18-0500

A compound of carbon,hydrogen and nitrogen is given and molecular mass of compound is given as 59

so gram molecular mass of compound =59g

now, as per the question it is given that , on combustion of the compound having mass .13 g it gives .29 g of "CO_2" and .18 g of "H_2O"

so, moles of compound goes on combustion reaction = "\frac{given mass}{molar mass}"

moles of compound on combustion = "\frac{.13}{59}"

moles of compound on combustion = .002203

moles of "CO_2 produced = \frac{.29}{44}"

moles of "CO_2" produced =.00659

moles of "H_2O" produced = "\frac{.18}{18}"

moles of "H_2O produced = .01"

now , from the above information we can conclude that

.002203 moles of compound gives .00659 moles of "CO_2"

so, one mole of compound gives "\frac{.00659}{.002203} = 3 moles of CO_2"

its means that the above compound has 3 atom of carbon

now, .002203 moles of compound gives .01 moles of water so

one mole of compound will given 4.55 mole of H2O

so , number of hydrogen atom ="2 \times 4.5 = 9"

now , it is given that

..145 g of compound gives 22.4 cm3 of N2 at STP

we know one mole has volume of 22400 cm3 of gas at STP

so moles of N2 produced = "\frac{22.4}{22400} = .001"

moles of compound gives this nitrogen = "\frac{.145}{59} = .00245"

one mole of compound gives N2 = 2.45

number of Nitrogen atom will be "\frac{2.45}{2} = 1.225" which is approx equal to 1

so here, we are getting 3 mole of C atom. 9 mole of Hydrogen atom and 1 mole of nitrogen atom


so its molecular formula is "C_3H_9N" it may be "C_3H_7 NH_2" its name may be Propanamine